I need a lot of help with theses problems. Your help will be greatly appreciated, thank you!
A. 3x - 6y = 12
B. 2x - y = -4
These two are the problems that I need help with, thank you a lot.

Answer by stanbon(48568)

These systems can be solved by graphing, substitution, and by elimination.
-------------------
A. 3x - 6y = 12
B. 2x - y = -4
----------------
Elimination:
Multiply equation B thru by 6 to get:
C. 12x -6y = -24
-------------------
Subtract A from C to get:
9x = -36
x = -4
----------
Substitute to solve for "y":
3(-4)-6y = 12
-12-6y = 12
6y = -24
y = -4
__________
Solution: (-4,-4)
====================
Cheers,
Stan H.
-